OpenStudy (cutiecomittee123):

HELP ME FILL IN THE BLANK FOR THIS EQUATION!! sin pi/4 sin pi/6 = 1/2 (cos pi/2 "blank" 5pi/12)

OpenStudy (cutiecomittee123):

I know that you can simplify some of the trig points using the unit circle sinpi/4=sqrt2/2 sinpi/6=1/2 cospi/2=0 so then (sqrt2/2)(1/2)=(1/2)(0) "blank" 5pi/12

OpenStudy (anonymous):

uh...

OpenStudy (cutiecomittee123):

\[(\sqrt2/2)(1/2)=(1/2)(0) "blank" 5\pi/12\]

OpenStudy (anonymous):

Im not sure...

OpenStudy (anonymous):

@malcolmmcswain @pooja195 @Jaynator495

OpenStudy (cutiecomittee123):

basically we need a trig function that can fit into the blank to make both sides equal, but that is where I am stumped!

OpenStudy (anonymous):

there is something wrong in your statement. right hand side is zero but left hand side is not equal to zero.

OpenStudy (baru):

she has left out the operation, +"blank"

OpenStudy (ac3):

\[\sin (\frac{ \pi }{ 4 })\sin(\frac{ \pi }{ 6 })=\frac{ 1 }{ 2 }(\cos (\frac{ \pi }{ 2 })blank(\frac{ 5\pi }{ 12 })\]

OpenStudy (ac3):

is this it?

OpenStudy (baru):

+blank otherwise rhs is zero

OpenStudy (ac3):

simplifying like you began is probably best

OpenStudy (ac3):

cospi/2 can't be right it makes the entire right side = 0 and the left side does not equal that.

OpenStudy (ac3):

doesn't matter what trig function you put in for blank that zero will make the right side equal zero no matter what you do

OpenStudy (baru):

shift evertything other than the blank to the left convert \(5\pi/12\) to \(\pi/4 +\pi/6\) use sin(A+B) formula and cos(A+B) formula and see what you get

OpenStudy (baru):

\[\sin (\frac{ \pi }{ 4 })\sin(\frac{ \pi }{ 6 })=\frac{ 1 }{ 2 }(\cos (\frac{ \pi }{ 2 })+blank(\frac{ 5\pi }{ 12 })\]

OpenStudy (ac3):

OHHHHHHH

OpenStudy (baru):

assume there is a plus before the blank, that is the only way its solvable

OpenStudy (ac3):

it's + blank

OpenStudy (ac3):

yea cuz otherwise it's an untrue statement

OpenStudy (baru):

\[\sin (\frac{ \pi }{ 4 })\sin(\frac{ \pi }{ 6 })=\frac{ 1 }{ 2 }(\cos (\frac{ \pi }{ 2 })blank(\frac{ 3 \pi +2\pi }{ 12 })\]

OpenStudy (ac3):

i'd help but she's not on to double check the equation

OpenStudy (baru):

\[\sin (\frac{ \pi }{ 4 })\sin(\frac{ \pi }{ 6 })=\frac{ 1 }{ 2 }(\cos (\frac{ \pi }{ 2 })+blank(\frac{ 3 \pi +2\pi }{ 12 })\]

OpenStudy (baru):

simplify and get blank(\(\frac{3\pi}{12}+\frac{2\pi}{12}\))=?

OpenStudy (baru):

@cutiecomittee123

OpenStudy (cutiecomittee123):

@baru sorrry I was cleaning my house up

OpenStudy (cutiecomittee123):

okay I took another look at the origional equaion and yes it is minus, not addition so it should be cos pi/2 - "blank" 5pi/12

OpenStudy (cutiecomittee123):

@Ac3 Hey I am back on, if you still want to help, you are totally welcome.

OpenStudy (baru):

we get blank(5pi/12)=\(-1/\sqrt{2}\) no trig function seems to solve this :(

OpenStudy (cutiecomittee123):

shoot... well there has to be a solution @baru

OpenStudy (cutiecomittee123):

technically the left side is equal to sqrt2/4 so the equation is sqrt2/4 = 1/2 (cos (pi/12) -____5pi/12)

OpenStudy (cutiecomittee123):

\[\sqrt2/4=1/2(\cos(\pi/12-____5\pi/12)\]

OpenStudy (baru):

hey, its cos(pi/12), everyone's been thinking its cos(pi/2), which is zero

OpenStudy (cutiecomittee123):

then our equation becomes \[\sqrt2/4=1/2((1+\sqrt3)/(2(\sqrt2))-____5\pi/12\]

OpenStudy (cutiecomittee123):

idk why it isnt presenting it correctly.

OpenStudy (cutiecomittee123):

but i am still stuck :((((

OpenStudy (cutiecomittee123):

@baru

OpenStudy (marcelie):

0

OpenStudy (cutiecomittee123):

what about 0?

OpenStudy (anonymous):

+3x

OpenStudy (anonymous):

3times

OpenStudy (anonymous):

1/12*(24k+-3/12)=1/12*(6 ? 5)

OpenStudy (anonymous):

k are integer

OpenStudy (cutiecomittee123):

wait where is the 24k from?

OpenStudy (anonymous):

cos(x)=sqrt(2)/2-> x= 2kpi+-pi/4

OpenStudy (cutiecomittee123):

okay I am not seeing how that works?

OpenStudy (baru):

Hey, can u use the equation Editor button and type out the question correctly.... Do double check, because it's unsolvable if u miss anything.... I have to leave now, but I'll have a go when I get back

OpenStudy (baru):

And should consider closing this thread and opening a new one....it's too cluttered and will scare off anyone else who might have an answer

OpenStudy (baru):

Tag me if u make a new one :) Cya

OpenStudy (anonymous):

\[R.H.S=\frac{ 1 }{ 2 }\left( \cos \frac{ \pi }{ 2 }Blank \frac{ 5 \pi }{ 12 } \right)=\frac{ 1 }{ 2 }\left( \cos \left( \frac{ \pi }{ 2 }\times \frac{ 6 }{ 5 \pi }\times \frac{ 5 \pi }{ 12 } \right) \right)\] \[=\frac{ 1 }{ 2 }\cos \frac{ \pi }{ 4 }=\frac{ 1 }{ 2 }*\frac{ 1 }{ \sqrt{2} }\]